# Improving Group Robustness on Spurious Correlation Requires Preciser Group Inference

### v1

@@ 3.2 GIC: Group Inference via data Comparison
As stated in Theorem 3.1, when $y^{c}$ is missing, we resort to maximizing $\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|\hat{y}^{c}_{s,\mathbf{w}}))$ as an alternative. We point out that maximizing a lower bound is meaningful as it provides the worst-case guarantee over the original objective. The detailed proof of Theorem 3.1 is provided in Appendix A.2.
@@
When implementing the objective (10) and (11), we further adopt a more computationally tractable form involving the difference between joint and marginal distributions to replace conditional probability distributions, e.g.,
$\begin{split}&\mathrm{KL}(\mathbb{P}(y^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}|\hat{y}^{c}_{s,\mathbf{w}}))=\mathrm{KL}(\mathbb{P}(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c},\hat{y}^{c}_{s,\mathbf{w}}))-\mathrm{KL}(\mathbb{P}(\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(\hat{y}^{c}_{s,\mathbf{w}})),\end{split}$ (12)
and we then employ the MINE algorithm [5] to estimate these $\mathrm{KL}(\cdot||\cdot)$ terms. The pseudocode, outlined in Algorithm 1, presents a comprehensive procedure for utilizing GIC to infer group labels.
Figure 1: Decision boundary visualization. $f_{\mathrm{ERM}}$ performs worse in recognizing spurious attributes than $f_{\mathrm{GIC}}$, and also worse in identifying invariant attributes than $f_{\mathrm{robust}}$.
Figure 2: Evaluation of group label inference. Compared to baselines, GIC significantly improves the recall while maintaining a relatively high precision.
@@ 3.3 Learning Robust Classifier with GIC
After obtaining the spurious attribute label $\hat{y}_{s}$, we then infer the group label as $\hat{g}=(y,\hat{y}_{s})$. The inferred groups $\hat{g}$ can be used with downstream invariant learning methods to learn invariant attributes and train a robust model $f_{\mathrm{robust}}$. For example, we can use Subsample technique to construct a balanced dataset by retaining all data from the smallest group and subsampling the data from the other inferred groups to match the same size for training $f_{\mathrm{robust}}$.
@@
In Section 4, we experiment with various downstream learning algorithms, demonstrating that the inferred groups from GIC can be flexibly utilized by diverse invariant learning algorithms to learn invariant features. This enables the training of robust models $f_{\mathrm{robust}}$ that efficiently mitigate spurious correlations.
3.4 How to Obtaion the Comparison Data
@@
Comparison data $\mathcal{C}$ with different group distributions from the training data is crucial in implementing GIC. The sources for comparison data are diverse, such as using a (labeled) validation data, or sampling from the unlabeled test data. In previous works, the validation set is often required to serve as a crucial basis for parameter selection [31, 51, 12], and even directly participate in model training, aiding in group inference [35].
@@
In challenging scenarios where both the validation set and the test set are inaccessible, we can manually create comparison data from the training set. This can be accomplished by resampling the training dataset in a non-uniform manner, such as adjusting the sampling weight based on the trained ERM’s prediction, as shown in JTT [31]. Specifically, the error set (missclassification set) of a trained ERM often represents minority groups where the spurious correlation no longer exists. By sampling fixed instances from the error set and non-error set, we can artificially construct comparison data with different group distributions.
@@
Although non-uniform sampling from the training set can ensure the availability of labeled comparison data, we still emphasize the reason we consider the unlabeled comparison data, such as sampling from the test set, is that in the real world, unlabeled data is often cheaper, abundant and easier to obtain [42, 18]. And subsequent analysis in Appendix C.4 supports that abundant comparison data is beneficial for the performance of GIC. By incorporating unlabeled comparison data, the applicability of GIC is enhanced.
Figure 3: Visualization of evaluated datasets with minority groups marked by red boxes. The spurious attribute and target exhibit strong spurious correlations, while these correlations typically does not hold for minority groups.
@@ 4 Experiments
Through our experimental evaluation, the primary goal is to address the following questions: (1) The effectiveness of GIC in mitigating spurious correlations: Can GIC successfully enhance worst-group accuracy and mitigate spurious correlation issues? (2) The accuracy of GIC in inferring group labels: Can GIC reliably predict group labels? (3) Analysis of misclassified cases: What factors contribute to misclassification by GIC for certain instances?
Our experimental results will be presented in the order of the three questions mentioned above. In the experiment, we use $\mathrm{GIC}_{\mathcal{C}_{y}}$ to denote the scenario where the comparison data has true labels, and $\mathrm{GIC}_{\mathcal{C}}$ to denote the scenario where the comparison data is unlabeled.
4.1 Experiments on Synthetic 2D Data
We start with a synthetic 2D dataset to demonstrate how GIC helps train a more robust model by learning spurious attributes. The synthetic dataset consists of training ($\mathcal{D}^{tr}$), validation ($\mathcal{D}^{val}$ which is the labeled comparison data $\mathcal{D}^{c}$), and test ($\mathcal{D}^{ts}$) sets. There is a spurious correlation in $\mathcal{D}^{tr}$, where true labels are highly correlated with the spurious attribute $\mathbf{x}_{1}$, while $\mathcal{D}^{c}$ and $\mathcal{D}^{ts}$ have correlations with the invariant feature $\mathbf{x}_{2}$. More details about the synthetic data can be found in Appendix B.1. Figure 2 shows the decision boundaries of the traditional ERM model $f_{\mathrm{ERM}}$, the spurious attribute classifier $f_{\mathrm{GIC}}$, and the retrained robust model $f_{\mathrm{robust}}$ using inferred group labels from $f_{\mathrm{GIC}}$ and the Subsample strategy. We also train ERM models exclusively on the invariant feature ($f_{\mathrm{invariant}}$) and the spurious feature ($f_{\mathrm{spurious}}$).
@@ 4.2 Experiments on Real-World Data
Real-World Datasets. We explore datasets in image and text classification that exhibit spurious correlations. For instance, CMNIST [2] involves digit recognition with spurious features where digit colors (red or green) are linked to digit values. Waterbirds [40] associates bird types with a spurious background attribute (water or land). CelebA [32] focuses on hair color recognition influenced by spurious gender-related features. CivilComments-WILDS [8, 23] aims to distinguish toxic from non-toxic online comments, with labels spuriously correlated with mentions of demographic identities. Figure 3 represents the spurious attributes and training targets of these datasets. Appendix B.2.1 provides a detailed group distribution description of these datasets. In the main text, we primarily use labeled and unlabeled validation sets as comparison data to demonstrate the effectiveness of GIC when comparison data is directly available. The experimental evidence on CMNIST in Appendix C.4 further illustrates the effectiveness of constructing comparison data through non-uniform sampling from the training set.
Baselines. For methods that address spurious correlation while requiring group labels, we consider GroupDRO [40], DFR [22], LISA [50], and SSA [35]. We also compare against methods that tackle spurious correlation without requiring group labels, namely, ERM, JTT [31], CnC [51], and EIIL [12]. ERM serves as a lower bound baseline, representing a basic training method without specific techniques to improve the accuracy of the worst-group. Additionally, we evaluate the accuracy of GIC in inferring group labels by comparing it to ERM and EI (the inferring group method of EIIL) and assess the GIC’s performance when using Mixup, Subsample, Upsample, and GrouDRO as downstream invariant learning methods. Detailed information on these methods can be found in Appendix B.2.2.
@@
Evaluation. In order to assess how well GIC tackles spurious correlations, we report the average-group and worst-group accuracies for all baselines. Furthermore, to evaluate the accuracy of GIC in inferring group labels, we report the precision (proportion of correctly inferred examples belonging to the true minority group) and recall (proportion of examples from the true minority group correctly inferred) of minority groups. We focus on minority groups because their small sample size presents challenges for accurate inference and successfully identifying these minority groups is crucial as spurious correlations often do not hold for them.
@@ 4.3 The effectiveness of GIC in mitigating spurious correlations
Table 1 showcases the average and worst-group accuracies for all methods. We specifically highlight GIC’s performance when combined with Mixup (denoted as $\mathrm{GIC}_{\mathcal{C}_{y}}$-M and $\mathrm{GIC}_{\mathcal{C}}$-M). In comparison to methods that train without leveraging group information, $\mathrm{GIC}_{\mathcal{C}_{y}}$-M and $\mathrm{GIC}_{\mathcal{C}}$-M consistently achieve higher worst-group accuracy across all datasets. Notably, even when compared to methods that incorporate group information during training, $\mathrm{GIC}_{\mathcal{C}_{y}}$-M and $\mathrm{GIC}_{\mathcal{C}}$-M deliver impressive results, particularly on CelebA and CivilComments datasets, where $\mathrm{GIC}_{\mathcal{C}_{y}}$-M almost matches the performance of baselines that utilize group labels. Furthermore, we observe that methods without group labels demonstrate weaker performance on worst-group accuracy compared to oracle group label-based methods. This discrepancy is especially pronounced when employing the same invariant learning algorithm (e.g., GroupDRO and EIIL, $\mathrm{GIC}_{\mathcal{C}_{y}}$-M and LISA). It underscores the necessity of further enhancing the accuracy of inferred group labels to boost the worst-group accuracy. The baselines, including JTT, CNC, EIIL, and SSA, tune all hyperparameters and employ early stopping based on the highest worst-group accuracy observed on the validation set. Similarly, we also consider using a validation set with group labels to ascertain the optimal number of training epochs. The ablation studies outlined in Appendix C.2 underscore the value of early stopping as a strategy within the framework of GIC.
Table 2: Worst group accuracy comparison (%). We highlight the best worst-group accuracy, mainly derived from GIC. The complete results with average-group accuracy are presented in Appendix C.1.
@@ 4.4 The accuracy of GIC in inferring group labels
We then delve deeper into comparing the performance of GIC, ERM, and EI in inferring group labels. In Figure 2, we find that GIC exhibits higher precision compared to the baselines on almost all datasets. This higher precision suggests that GIC’s estimated minority groups have low redundancy, thereby increasing the likelihood of sampling true minority group examples to assist in training robust models. Furthermore, GIC’s primary advantage lies in its high recall, consistently maintained at over 60% and even surpassing 80% in CMNIST, CelebA, and CivilComments. The high recall rate indicates the estimated minority groups by GIC contain more diverse samples from true minority groups which can provide more discriminative information for downstream invariant learning tasks [17].
In Table 3, we present our findings on combining GIC, ERM, and EI with various invariant learning algorithms, including GroupDRO [13], Subsample [22], and Upsample [31]. Our worst-group results show that GIC consistently outperforms the baselines across different downstream algorithms, further emphasizing its superiority in group inference and improving worst-group performance. The ablation results are also shown in Appendix C.2.
@@ 4.5 Error Case Analysis
We then visualize misclassified samples and focus on the underlying factors contributing to these errors. Figure 4 shows examples from the CelebA dataset, where the spurious attribute is gender (woman and man), and the association between hair color and gender is considered spurious,such as blonde hair is correlated with women. We notice an interesting phenomenon called semantic consistency in GIC. For example, GIC misclassifies women with short hair as men, who have a strong semantic resemblance to correctly classified (short hair) male samples. We also observe a similar phenomenon in the waterbirds dataset in Appendix 4.5, where GIC tends to classify water backgrounds with prominent land elements (such as trees) as land backgrounds. This semantic consistency significantly contributes to prediction errors in GIC and has a negative effect on downstream algorithms that rely on sampling. It can lead to an overrepresentation of majority group samples (e.g., [blond, women]) within the estimated minority group ([blond, men]), exacerbating group imbalances during resampling.
However, for methods like Mixup [50] that aim to disrupt spurious correlations between spurious attributes and true labels for invariant learning, high semantic consistency is meaningful. We then show how GIC’s semantic consistency leads to better mixed images compared to using oracle group labels. In Figure 5, Mixup disrupts the spurious correlation between blonde hair and women by mixing samples from the same class (blond) but with different spurious attributes (man). When sampling from the oracle blond man group, long-haired men may be selected, resulting in mixed images that still retain the typical woman attribute (long hair). However, by using the blond man sample inferred by GIC, such as the short-haired woman in Figure 5, the generated mixed samples is more closely resemble blonde men compared to using oracle group labels. These findings may explain the better performance of $\mathrm{GIC}_{\mathcal{C}_{y}}$-M than LISA (which uses oracle group labels with Mixup) in Table 1, even though GIC’s precision and recall are not at 100%.
Figure 5: GIC generates better mixed images. By leveraging the high semantic consistency in image recognition, spurious attributes and true labels are decoupled in mixed images generated by GIC.

### v2

@@ 3.2 GIC: Group Inference via data Comparison
In Theorem 3.1, when $y^{c}$ is missing, we resort to maximizing the lower bound $\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|\hat{y}^{c}_{s,\mathbf{w}}))$ as an alternative. We point out that maximizing a lower bound is meaningful as it provides the worst-case guarantee over the original objective. The detailed proof of Theorem 3.1 is provided in Appendix A.1.
@@
When implementing the objective (11) and (12), we further adopt a more computationally tractable form involving the difference between joint and marginal distributions to replace conditional probability distributions. Taking the case with the label $y^{c}$ as an example, we consider
$\begin{split}&\mathrm{KL}(\mathbb{P}(y^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}|\hat{y}^{c}_{s,\mathbf{w}}))\\ &=\mathrm{KL}(\mathbb{P}(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c},\hat{y}^{c}_{s,\mathbf{w}}))-\mathrm{KL}(\mathbb{P}(\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(\hat{y}^{c}_{s,\mathbf{w}})),\end{split}$ (13)
and then the MINE algorithm (Belghazi et al., 2018) is employed to estimate these $\mathrm{KL}(\cdot||\cdot)$ terms. The pseudocode, outlined in Algorithm 1, presents a comprehensive procedure for utilizing GIC to infer group labels.
Figure 3: Visualization of evaluated datasets with minority groups marked by red boxes. The spurious attribute and targets exhibit strong spurious correlations, while these correlations typically does not hold for minority groups.
@@ 3.3 Learning Robust Classifier with GIC
After obtaining the spurious attribute label $\hat{y}_{s,\mathbf{w}}$, we then infer the group label as $\hat{g}=(y,\hat{y}_{s,\mathbf{w}})$. The inferred groups $\hat{g}$ can be used with downstream invariant learning methods to learn invariant attributes and train a robust model $f_{\mathrm{robust}}$. For example, we can use Subsample technique to construct a balanced dataset by retaining all data from the smallest group and subsampling the data from the other inferred groups to match the same size for training $f_{\mathrm{robust}}$.
@@
In Section 4, we experiment with various downstream learning algorithms, demonstrating that the inferred groups from GIC can be flexibly utilized by diverse invariant learning algorithms to learn invariant features. This facilitates the training of robust models, denoted by $f_{\mathrm{robust}}$, which effectively mitigate spurious correlations.
3.4 How to Obtain the Comparison Data
@@
Comparison data $\mathcal{C}$ with different group distributions from the training data is crucial in implementing GIC. The sources for comparison data are diverse, such as using a (labeled) validation data, or sampling from the unlabeled test data. In previous works, the validation set is often required to serve as a crucial basis for parameter selection (Liu et al., 2021; Zhang et al., 2022; Creager et al., 2021), and even directly participate in model training, aiding in group inference (Nam et al., 2022).
@@
In challenging scenarios where both the validation set and the test set are inaccessible, we can manually create comparison data from the training set. This can be accomplished by resampling the training dataset in a non-uniform manner, such as adjusting the sampling weight based on the trained ERM’s prediction, as shown in JTT (Liu et al., 2021). Specifically, the error set (missclassification set) of a trained ERM often represents minority groups where the spurious correlation no longer exists. By sampling from the error set and non-error set, we can artificially construct comparison data with different group distributions.
@@
Although non-uniform sampling from the training set can ensure the availability of labeled comparison data, we still emphasize the reason we consider the unlabeled comparison data, such as sampling from the test set, is that in the real world, unlabeled data is often cheaper, abundant and easier to obtain (Shejwalkar et al., 2023; Göpfert et al., 2019). And subsequent analysis in Appendix C.4 supports that abundant comparison data is beneficial for the performance of GIC. By incorporating unlabeled comparison data, the applicability of GIC is enhanced.
3.5 The Relationship between GIC and ERM
In this section, we highlight the relationship between GIC and ERM. When there is no difference in group distribution between training and comparison data, or the weighting parameter $\gamma=0$, GIC degenerates to ERM in terms of group label inference. the KL loss in (11) and (12) is minimal, indicating a reduced impact of the spurious term. Consequently, the CE loss becomes predominant, causing GIC to effectively become an ERM-based group inference method.
Although ERM can still be used for inferring group labels, it lacks the crucial spurious term necessary to violate the principle of invariant learning and to effectively identify spurious features. Thus, ERM-based inference should serve as the performance baseline for GIC’s group label performance which is validated by subsequent experiments. Moreover, GIC can also be viewed as a special ERM-based group inference method where the spurious term functions as a regularization term. By incorporating insights from the comparison data, the spurious term encourages the trained neural network to differentiate between spurious and invariant features, thus enhancing group label inference.
@@ 4 Experiments
Through our experimental evaluation, the primary goal is to address the following questions: (1) The effectiveness of GIC in mitigating spurious correlations: Can GIC successfully enhance worst-group accuracy and mitigate spurious correlation issues? (2) The accuracy of GIC in inferring group labels: Can GIC reliably predict group labels? (3) Analysis of misclassified cases: What factors contribute to misclassifications by GIC for certain instances?
Our experimental results will be presented in the order of the three questions mentioned above. In experiments, we use $\mathrm{GIC}_{\mathcal{C}_{y}}$ to denote the scenario where the comparison data has true labels, and $\mathrm{GIC}_{\mathcal{C}}$ to denote the scenario where the comparison data is unlabeled.
@@
4.1 Experiments on Synthetic 2D Data
We start with a synthetic 2D dataset to demonstrate how GIC helps train a more robust model by learning spurious attributes. The synthetic dataset consists of training ($\mathcal{D}^{tr}$), validation ($\mathcal{D}^{val}$ which is the labeled comparison data $\mathcal{D}^{c}$), and test ($\mathcal{D}^{ts}$) sets. There is a spurious correlation in $\mathcal{D}^{tr}$, where true labels are highly correlated with the spurious attribute $\mathbf{x}_{1}$, while $\mathcal{D}^{c}$ and $\mathcal{D}^{ts}$ have correlations with the invariant feature $\mathbf{x}_{2}$. More details about the synthetic data can be found in Appendix B.1. Figure 2 shows the decision boundaries of the traditional ERM model $f_{\mathrm{ERM}}$, the spurious attribute classifier $f_{\mathrm{GIC}}$, and the retrained robust model $f_{\mathrm{robust}}$ using inferred group labels from $f_{\mathrm{GIC}}$ and the Subsample strategy. We also train ERM models exclusively on the invariant feature (denoted as $f_{\mathrm{invariant}}$) and the spurious feature (denoted as $f_{\mathrm{spurious}}$).
@@ 4.2 Experiments on Real-World Data
Real-World Datasets. We explore datasets in image and text classification that exhibit spurious correlations. For instance, CMNIST (Arjovsky et al., 2019) involves digit recognition with spurious features where digit colors (red or green) are linked to digit values. Waterbirds (Sagawa et al., 2019) associates bird types with a spurious background attribute (water or land). CelebA (Liu et al., 2015) focuses on hair color recognition influenced by spurious gender-related features. CivilComments-WILDS (Borkan et al., 2019; Koh et al., 2021) aims to distinguish toxic from non-toxic online comments, with labels spuriously correlated with mentions of demographic identities. Figure 3 represents the spurious attributes and training targets of these datasets. Appendix B.2.1 provides a detailed group distribution description of these datasets. In the main text, we primarily use labeled and unlabeled validation sets as comparison data to demonstrate the effectiveness of GIC when comparison data is directly available. The experimental evidence in Appendix C.4 further illustrates the effectiveness of constructing comparison data through non-uniform sampling from the training set.
Baselines. For methods that address spurious correlation while requiring group labels, we consider GroupDRO (Sagawa et al., 2019), DFR (Kirichenko et al., 2022), LISA (Yao et al., 2022), and SSA (Nam et al., 2022). We also compare against methods that tackle spurious correlation without requiring group labels, namely, ERM, JTT (Liu et al., 2021), CnC (Zhang et al., 2022), and EIIL (Creager et al., 2021). ERM serves as a lower bound baseline, representing a basic training method without specific techniques to improve the accuracy of the worst-group. Additionally, we evaluate the accuracy of GIC in inferring group labels by comparing it to ERM and EI (the inferring group method of EIIL) and assess the GIC’s performance when using Mixup, Subsample, Upsample, and GrouDRO as downstream invariant learning methods. Detailed information on these methods can be found in Appendix B.2.2.
@@
Evaluation. In order to assess how well GIC tackles spurious correlations, we report the average and worst-group accuracy for all baselines. Furthermore, to evaluate the accuracy of GIC in inferring group labels, we report the precision (proportion of correctly inferred examples belonging to the true minority group) and recall (proportion of examples from the true minority group correctly inferred) of minority groups. We focus on minority groups because their small sample size presents challenges for accurate group inference and successfully identifying these minority groups is crucial as spurious correlations often do not hold for them.
@@ 4.3 The Effectiveness of GIC in Mitigating Spurious Correlations
Table 2: Worst-group accuracy comparison (%). We highlight the 1st worst-group accuracy, mainly derived from GIC. The complete results including average-group accuracy are presented in Appendix C.1.
@@
Table 1 showcases the average and worst-group accuracies for all methods. We specifically highlight GIC’s performance when combined with Mixup (denoted as $\mathrm{GIC}_{\mathcal{C}_{y}}$-M and $\mathrm{GIC}_{\mathcal{C}}$-M). In comparison to methods that train without leveraging group information, $\mathrm{GIC}_{\mathcal{C}_{y}}$-M and $\mathrm{GIC}_{\mathcal{C}}$-M consistently achieve higher worst-group accuracy across all datasets. Notably, even when compared to methods that incorporate group information during training, $\mathrm{GIC}_{\mathcal{C}_{y}}$-M and $\mathrm{GIC}_{\mathcal{C}}$-M deliver impressive results, particularly on CelebA and CivilComments datasets, where $\mathrm{GIC}_{\mathcal{C}_{y}}$-M almost matches the performance of baselines that utilize group labels. Furthermore, we observe that methods without group labels demonstrate weaker performance on worst-group accuracy compared to oracle group label-based methods. This discrepancy is especially pronounced when employing the same invariant learning algorithm (e.g., GroupDRO and EIIL, $\mathrm{GIC}_{\mathcal{C}_{y}}$-M and LISA). It underscores the necessity of further enhancing the accuracy of inferred group labels to boost the worst-group accuracy. The baselines, including JTT, CNC, EIIL, and SSA, tune hyperparameters and employ early stopping based on the highest worst-group accuracy observed on the validation set. Similarly, we also consider using the highest worst-group accuracy of the validation set with group labels as a criterion to ascertain the optimal number of training epochs. The ablation studies detailed in Appendix C.2 highlight the importance of early stopping as a strategy in the GIC framework.
@@ 4.4 The Accuracy of GIC in Inferring Group Labels
We then delve deeper into comparing the performance of GIC, ERM, and EI in inferring group labels of minority groups. In Figure 2, we find that GIC exhibits higher precision compared to the baselines on almost all datasets. This higher precision suggests that the minority groups estimated by GIC have low redundancy, thereby increasing the possibility of utilizing true minority group examples to assist in training robust models for downstream tasks. Furthermore, GIC’s primary advantage lies in its high recall, consistently maintained at over 60% and even surpassing 80% in CMNIST, CelebA, and CivilComments. The high recall rate indicates the estimated minority groups by GIC contain more diverse samples from true minority groups which can provide more discriminative information for downstream invariant learning tasks (Gong et al., 2019).
In Table 3, we present our findings on combining GIC, ERM, and EI with various invariant learning algorithms, including GroupDRO (Duchi et al., 2019), Subsample (Kirichenko et al., 2022), and Upsample (Liu et al., 2021). Our worst-group results show that GIC consistently outperforms the baselines across different downstream algorithms, further emphasizing its superiority in group inference and worst-group performance improvement. The ablation results concerning the use of the early stopping strategy when training robust models are also shown in the Appendix C.2.
@@ 4.5 Error Case Analysis
In this section, we then focus on the underlying factors contributing to these errors via visualizing the misclassified samples. Figure 4 shows examples from the CelebA dataset, where the spurious attribute is gender (woman and man), and the association between hair color and gender is considered spurious, such as blonde hair is correlated with women. We note an interesting phenomenon known as semantic consistency in GIC. For instance, GIC misclassifies women with short hair as men, who bear a strong semantic resemblance to correctly classified male samples with short hair. This phenomenon is also evident in the waterbirds dataset, as detailed in Appendix 4.5, where GIC tends to classify water backgrounds with prominent land elements (such as trees) as land backgrounds. This semantic consistency significantly contributes to prediction errors in GIC and has a negative effect on downstream algorithms that rely on sampling. It can lead to an overrepresentation of majority group samples ([blond, women]) within the estimated minority group ([blond, men]), exacerbating group imbalances during resampling.
Figure 5: GIC generates better mixed images. By leveraging the high semantic consistency in image recognition, spurious attributes and true labels are decoupled in mixed images generated by GIC. Various mixing techniques are employed to handle different datasets, as detailed in Section B.2.2.
However, for invariant methods like Mixup (Yao et al., 2022) that aim to disrupt spurious correlations between spurious attributes and true labels for invariant learning, high semantic consistency is beneficial. We then show how GIC’s semantic consistency leads to better mixed images compared to using oracle group labels. In Figure 5, Mixup disrupts the spurious correlation between blonde hair and women by mixing samples from the same class (blond) but with different spurious attributes (man). When sampling from the oracle blond man group, long-haired men may be selected, resulting in mixed images that still retain the typical woman attribute (long hair). However, by using the blond man sample inferred by GIC, such as the short-haired woman in Figure 5, the generated mixed samples is more closely resemble blonde men compared to using oracle group labels. Similarly, advantages can also occur in the waterbirds dataset as shown in Figure 5. These findings may explain the better performance of $\mathrm{GIC}_{\mathcal{C}_{y}}$-M than LISA (which uses oracle group labels with Mixup) in Table 1, even though GIC’s precision and recall are not at 100%.